OpenStudy (anonymous):

check wheather the ordered pairs is a soultion of the system first one to answer gets a medal (-1,1) x+y=0 3x-3y=0

OpenStudy (lalaly):

-1+1 = 0 -3-3=-6 so not a solution

OpenStudy (anonymous):

no for both or just for one?